OpenStudy (anonymous):

There were twice as many carnations is roses in the flower shop. After selling 50 carnations and 10 roses, there were three times as many roses carnations left in the shop. How many roses were there in the shop at first?

OpenStudy (anonymous):

A Mathematica calculation is shown below: \[\text{Eliminate}\left[\left\{\frac{c}{2}=r,\frac{c-50}{r-10}=\frac{1}{3}\right\},c\right] \]r = 28